The number of degrees of freedom of a rigid body in n- dimensional space is
1
Expert's answer
2019-04-15T10:00:14-0400

A rigid body in n- dimensional space has n translational degrees of freedom and


"\frac{n(n-1)}{2}"

rotational degrees of freedom. 

The number of degrees of freedom of a rigid body in n- dimensional space is


"n+\frac{n(n-1)}{2}=\frac{n(n+1)}{2}"